OpenStudy (ilovemath14):

1. (6 to the 2 power) to the 2 power 2. (5 to the 2 power) to the 3 power 3. (-5q to the 2 power p) to the 3 power

OpenStudy (anonymous):

1) \[(6^2)^2\] 2) \[(5^2)^3\] 3) \[((-5q)^{2p})^3\]

OpenStudy (anonymous):

use this property \[(a^b)^c=a^{bc}\]

OpenStudy (ilovemath14):

This is what I need to answer

OpenStudy (anonymous):

ok \[(6^2)^2=6^{2*2}=6^4\]

OpenStudy (anonymous):

1296 is 6^4

OpenStudy (anonymous):

\[(5^2)^3=5^{2*3}=5^6=15625\]